It says to evaluate(solve) the arithmetic series:
20+27+34+41..., n=16
How do you do this? Please help. Thanks!

Answer by rothauserc(4718)

The sum of an arithmetic sequence is given by the following formula,
(n/2)*(a1 + an), which is the sum of the first n terms
the nth term an is given by the following formula
an = a1 + (n-1)d, where d is the common difference, therefore
a16 = 20 + (16-1)*7
a16 = 20 + 105 = 125
now the sum is (16/2)*(20 + 125) = 1160